# Twisted Superfields

### v8

@@ 5.1 A Light Z’ Boson
In section 4, it is mentioned that the top quark generates an anomaly for $Z^{\prime}$. This is actually assumed to first affect $Y^{\prime}$ from the top quark scale to the $Z$ boson mass scale, and then to affect $Z^{\prime}$ below the $Z$ boson mass scale. That anomaly causes the coupling for $Y^{\prime}$ interactions with $Q_{21}$ and gaugino fields (both of which have top-quark components) to run down more quickly than the coupling for $Y^{\prime}$ interactions with $Q_{22}$ fields (which only have bottom and down quark components). That effect can be parametrized by a factor $\kappa_{22}$ by which the $g_{Y^{\prime}}$ coupling to $Q_{22}$ exceeds that to $Q_{21}$ and gaugino fields at the scale of $m_{Z}$.
In that case, then the $T^{Z^{\prime}}_{22}$ group structure is derived by the following modification of the Weinberg rotation of eq (2.16):
$g_{Z^{\prime}}T^{Z^{\prime}}_{22}=$ $\kappa_{22}\cos\phi_{Z}g_{Z^{\prime}}T^{Z^{\prime}}_{22}$ (5.1)
@@
$g_{Z^{\prime}}T_{G}^{Z^{\prime}}$ $={\rm{diag}}\left(\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,\tfrac{1}{2}+\tfrac{1}{3}x,x-1,0,0\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}$
$g_{Z^{\prime}}T_{22}^{Z^{\prime}}$ $=\tfrac{1}{2}\kappa^{\prime}_{22}\left(\kappa_{22}{\rm{diag}}(1,1,1,-1,-1,0)+\left(\tfrac{2}{3}x,\tfrac{2}{3}x,\tfrac{2}{3}x,2x-1,1,0\right)\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}\,,$
where $T_{G}^{Z^{\prime}}$ is the group structure when acting on gauginos, and $\kappa^{\prime}_{22}$ is a second parameter that expresses the fact that below $m_{Z}$ but above $m_{b}$, the $g_{Z^{\prime}}$ coupling to $Q_{22}$ runs down more slowly than the $g_{Z^{\prime}}$ coupling to $Q_{21}$ and gaugino fields. The $Z^{\prime}$ group structure when interacting with adjoint fields is assumed to be the same as $T_{22}^{Z^{\prime}}$, since adjoint fields also do not have any top-quark components.
@@
$\begin{array}[]{|r|c|c|}\hline\cr&g_{L}&g_{R}\\ \hline\cr u&-1.15&-0.65\\ d&3.0&0.71\\ e&-0.15&-0.15\\ \nu_{e}&0&0\\ \hline\cr\end{array}\,\,\,\times\frac{e\sin\phi_{Z}}{\sin\theta_{W}\cos\theta_{W}}\,.$ (5.3)
It was assumed for the fit that $\tfrac{1}{2}\kappa^{\prime}_{22}\left(\kappa_{22}+\tfrac{2}{3}x\right)\simeq 3.0$.
@@
$\begin{array}[]{|r|c|c|}\hline\cr&V&A\\ \hline\cr p&C_{p}\simeq 0.02&a_{p}\simeq 0.25\\ n&C_{n}\simeq 1.1&a_{n}\simeq-0.44\\ e&|\epsilon|\simeq 1.0&0\\ \hline\cr\end{array}\,\,\,\times 10^{-3}\,,$ (5.5)
@@
where notation from [33] (Appendix C) is being used for protons and neutrons, and $\epsilon$ is the electron vector coupling divided by $e\simeq 0.3$.
@@
With these choices, the $Z^{\prime}$ boson of this model can reproduce $X17$ data while satisfying all applicable constraints. From fig. 5 of [33], it can be seen that the proton and neutron vector couplings can reproduce the $X17$ 12C data, while the axial vector couplings can reproduce the 8Be and 4He data. The vector coupling is protophobic enough to satisfy the NA48 constraint. The fact that there are no interactions with electron neutrinos allows the model to satisfy neutrino constraints [34]. The electron interaction strength is above the limit imposed by NA64 but below the level that would contradict electron $g-2$ data [35]. Finally, as described in Appendix F of [33], the very stringent atomic parity constraints on mixed-parity models do not apply when the electron has only a vector interaction (no axial vector), as in this model.

### v9

@@ 5.10 A Light Z’ Boson
In section 4, it is mentioned that the heavy leptons generate an anomaly for $Z^{\prime}$. Below the mass of the top quark, the $Z^{\prime}$ anomaly is exacerbated when the top quark decouples, since the anomaly is not resolved by the $Z$ boson mass. Below the top-quark mass scale, it is assumed that the coupling for $Y^{\prime}$ interactions with $Q_{22}$ fields (with an effective quark anomaly) run down more quickly than those with $Q_{21}$ fields (without one). This effect is parametrized by a factor $\kappa$ by which the $g_{Y^{\prime}}$ coupling to $Q_{22}$ is less than its coupling to $Q_{21}$ at the $Z^{\prime}$ mass scale.
The $T^{Z^{\prime}}_{22}$ group structure is then derived by the following modification of the Weinberg rotation of eq (2.16):
$g_{Z^{\prime}}T^{Z^{\prime}}_{22}=$ $\kappa\cos\phi_{Z}g_{Y^{\prime}}T^{Y^{\prime}}_{22}$ (5.15)
@@
$g_{Z^{\prime}}T_{22}^{Z^{\prime}}$ $={\rm{diag}}\left(\tfrac{1}{2}\kappa+\tfrac{1}{3}x,\tfrac{1}{2}\kappa+\tfrac{1}{3}x,\tfrac{1}{2}\kappa+\tfrac{1}{3}x,x-\tfrac{1}{2}-\tfrac{1}{2}\kappa,\tfrac{1}{2}-\tfrac{1}{2}\kappa,0\right)e\sin\phi_{Z}/\sin\theta_{W}\cos\theta_{W}\,$,
where $x=\sin^{2}\theta_{W}$. The $Z^{\prime}$ group structure when interacting with gaugino and adjoint fields is the same as $T_{22}^{Z^{\prime}}$.
@@
$\begin{array}[]{|r|c|c|}\hline\cr&g_{L}&g_{R}\\ \hline\cr u&1+\tfrac{2}{3}x&-\tfrac{1}{2}-\kappa+\tfrac{2}{3}x\\ d&\tfrac{1}{2}\kappa+\tfrac{1}{3}x&\tfrac{1}{2}\kappa+\tfrac{1}{3}x\\ e&x-1&x-1\\ \nu_{e}&0&0\\ \hline\cr\end{array}\,\,\,\times\frac{e\sin\phi_{Z}}{\sin\theta_{W}\cos\theta_{W}}\,.$ (5.17)
In order to make the $Z^{\prime}$ protophobic (see below), the following phenomenological choice is made:
$\kappa\simeq\tfrac{1}{3}+\tfrac{2}{3}x\,.$ (5.18)
@@
$\begin{array}[]{|r|c|c|}\hline\cr&V&A\\ \hline\cr p&C_{p}\simeq 0&a_{p}\simeq 5.4\\ n&C_{n}\simeq 1.5&a_{n}\simeq-1.1\\ e&|\epsilon|\simeq 7.6&0\\ \hline\cr\end{array}\,\,\,\times 10^{-4}\,$, (5.20)
@@
where notation from [71] (Appendix C) is being used for protons and neutrons, $x\simeq.238$ was used, and $\epsilon$ is the electron vector coupling divided by $e\simeq 0.3$.
@@
With these choices, the $Z^{\prime}$ boson of this model can reproduce $X17$ data while satisfying all applicable constraints. From fig. 5 of [71], it can be seen that the proton and neutron vector couplings can reproduce the $X17$ 12C data, while the axial vector couplings can reproduce the 8Be and 4He data. The vector coupling is protophobic ($C_{p}=0$), satisfying the NA48 constraint. The fact that there are no interactions with electron neutrinos allows the model to satisfy neutrino constraints [72]. The electron interaction strength is above the limit imposed by NA64 but below the level that would contradict electron g2g-2 data [73]. Finally, as described in Appendix F of [71], the very stringent atomic parity constraints on mixed-parity models do not apply when the electron has only a vector interaction (no axial vector), as in this model.